# Illuminating gravitational wave sources with Sgr A* flares

### v1

@@ Abstract
Sagittarius A* (Sgr A*), the supermassive black hole at the center of the Milky Way, exhibits daily energetic flares characterized by non-thermal emission in the infrared and X-ray bands. While the underlying accretion flow is a Radiatively Inefficient Accretion Flow (RIAF) peaking at radio frequencies, the mechanism powering these non-thermal transients remains debated. Stellar dynamics predict a population of faint brown dwarfs orbiting Sgr A*. These objects are progenitors of Extremely Large Mass Ratio Inspirals (XMRIs), crucial sources of low-frequency gravitational waves for the future Laser Interferometer Space Antenna (LISA) mission. We investigate whether the tidal stripping of brown dwarfs provides a viable fueling mechanism for the observed flares. Here we present high-resolution hydrodynamic simulations of grazing tidal interactions coupled with a parameterized non-thermal radiation model. We demonstrate that the dynamics of the tidal fallback and subsequent viscous evolution naturally reproduce the fundamental temporal characteristics of observed flares: the peak luminosity and the characteristic 1-hour duration. We show that this fueling mechanism is dynamically viable and energetically consistent, placing strong constraints on the required efficiency of the non-thermal emission process, suggesting extreme radiative inefficiency. These findings provide compelling evidence for a hidden population of brown dwarfs in the Galactic Center. Crucially, the observed high flare frequency implies tight orbits characteristic of advanced inspirals. This establishes a direct link between electromagnetic transients and active gravitational wave sources, alerting the LISA consortium years in advance to the presence of specific XMRI systems promising exceptionally high signal-to-noise ratios for precision tests of general relativity.
Sagittarius A* (1413) — Galactic center (565) — Supermassive black holes (1663) — Brown dwarfs (185) — Gravitational wave sources (677)
software: MESA (B. Paxton et al., 2011, 2013; P. P. Eggleton, 1983), StarSmasher (F. A. Rasio, 1991; J. C. Lombardi, 1998; F. A. Rasio & J. C. Lombardi, 1999; J. C. Lombardi et al., 1999)
show]amaro@upv.es
@@ 1 Introduction
While the emission mechanism is identified as non-thermal synchrotron radiation, the mechanism triggering these events—specifically, the source of the fuel and the trigger for particle acceleration—remains debated. Proposed fueling mechanisms include the interaction with the G2 object, accretion instabilities driven by stellar winds, or large-scale magnetic activity  (O. Pfuhl et al., 2015; G. Witzel et al., 2014; G. Ponti et al., 2015; D. Calderón et al., 2025; S. M. Ressler et al., 2019).
We investigate the hypothesis that these flares are fueled by the tidal stripping of material from a brown dwarf (BD) during a grazing encounter with Sgr A*. Stellar dynamics models predict a substantial population of BDs in the Galactic Center. The interaction of these objects with the supermassive black hole has implications beyond electromagnetic flares. BDs inspiraling toward Sgr A* form Extremely-Large Mass Ratio Inspirals (XMRIs), characterized by mass ratios of $\sim 10^{8}$. These systems are key sources of low-frequency gravitational waves detectable by the future Laser Interferometer Space Antenna (LISA) mission (P. Amaro-Seoane, 2019; P. Amaro Seoane & S.-D. Zhao, 2025). Modeling the electromagnetic signatures of these tidal encounters provides a means to constrain the population of these otherwise undetectable objects.
@@
2 Methodology
This study utilizes high-resolution hydrodynamic simulations coupled with a parameterized radiative post-processing model. This section details the construction of the progenitor model, the simulation techniques, the physical constraints leading to the simulation strategy, and the analysis methods employed.
2.1 Brown Dwarf Model Construction
@@
We constructed realistic initial conditions for the progenitor brown dwarf (BD) using the Modules for Experiments in Stellar Astrophysics (MESA) code (P. P. Eggleton, 1983; B. Paxton et al., 2011, 2013). We modeled the progenitor as a $0.01\,M_{\odot}$ object evolved for a Hubble time. The resulting stabilized one-dimensional structure is characterized by a logarithmic radius $\log_{10}(R/R_{\odot})=-0.9833$ and a logarithmic effective temperature $\log_{10}(T_{\text{eff}}/\text{K})=2.3447$. The composition is defined by mass fractions of hydrogen $X=0.7597$, helium $Y=0.2402$, and metals $Z=10^{-4}$.
@@
2.2 Hydrodynamic Simulations
We performed the hydrodynamic simulations of the tidal encounter using the StarSmasher SPH code (F. A. Rasio, 1991; J. C. Lombardi, 1998; F. A. Rasio & J. C. Lombardi, 1999; J. C. Lombardi et al., 1999). The stabilized BD model is placed on specified Keplerian orbits characteristic of XMRI progenitors around Sgr A*, which is modeled as a $4.3\times 10^{6}\,M_{\odot}$ point mass sink.
StarSmasher employs variational equations of motion to evolve the system. It calculates hydrodynamic forces by determining the pressure of each particle as a function of its local properties via the equation of state. To calculate gravitational interactions, the code implements a direct $N$-body summation method accelerated on NVIDIA graphics processing units (GPUs). StarSmasher utilizes a cubic spline function for the smoothing kernel. To manage shocks and prevent unphysical interparticle penetration, the code employs an artificial viscosity prescription coupled with a Balsara Switch.
2.3 Tidal Interaction and Renormalization Strategy
We analyze the dynamics of the interaction and the constraints imposed by SPH resolution, which necessitates the use of high-mass proxy simulations followed by renormalization.
2.3.1 Analytical Estimates of Tidal Stripping
We first estimate the parameters of the tidal encounter for a representative orbit. We consider the $0.01\,M_{\odot}$ BD orbiting Sgr A* ($M_{\text{BH}}=4.3\times 10^{6}\,M_{\odot}$) with an illustrative periastron distance $R_{p}=120\,R_{\odot}$ and semi-major axis $a=2560\,R_{\odot}$. Using the derived BD radius $R_{\text{BD}}\approx 0.1\,R_{\odot}$, the tidal radius is
@@ 2.3.1 Analytical Estimates of Tidal Stripping
For such encounters, analytical models predict the fraction of stellar mass stripped. We estimate the stripped mass using the formalism of tidal energy deposition. For a penetration factor $\beta$, the energy injected into the BD structure relative to its binding energy is given by
@@
Adopting a representative tidal coupling coefficient $T(\beta)\sim 10^{-4}$ appropriate for our scenario, we find $\Delta E_{\text{tide}}/E_{\text{bind}}\approx 6.3\times 10^{-6}$. For an $n=1.5$ polytropic structure appropriate for a BD, this yields an estimate of the analytically expected stripped mass:
$\Delta M_{\text{analytical}}\approx 3.8\times 10^{-6}\,M_{\odot}.$ (3)
2.3.2 Resolution Constraints and Proxy Simulations
We now examine the resolution constraints of our SPH simulations. The mass resolution of the simulation, defined by the mass of a single SPH particle, $m_{\text{SPH}}$, is:
@@ 2.3.2 Resolution Constraints and Proxy Simulations
The critical constraint arises when considering the physical mass required to power the observed flare. As determined by the optimization detailed in Section 3, the required physical accreted mass is $M_{\text{physical}}\approx 1.5\times 10^{-7}\,M_{\odot}$ (assuming a derived non-thermal radiative efficiency $\eta_{\text{NT}}\approx 10^{-8}$).

### v2

@@ Abstract
Context. Sagittarius A* (Sgr A*), the supermassive black hole at the center of the Milky Way, exhibits daily energetic flares characterized by non-thermal emission in the infrared and X-ray bands. While the underlying accretion flow is a Radiatively Inefficient Accretion Flow (RIAF) peaking at radio frequencies, the mechanism powering these non-thermal transients remains debated. Stellar dynamics predict a population of faint brown dwarfs orbiting Sgr A*.
Aims. We investigate whether the tidal stripping of brown dwarfs provides a viable fueling mechanism for the observed flares. These objects are progenitors of Extremely Large Mass Ratio Inspirals (XMRIs), crucial sources of low-frequency gravitational waves for the future Laser Interferometer Space Antenna (LISA) mission.
Methods. We present high-resolution hydrodynamic simulations of grazing tidal interactions coupled with a parameterized non-thermal radiation model. We utilize Smoothed Particle Hydrodynamics (SPH) to model the stripping of the brown dwarf envelope and the subsequent accretion of this material. Due to computational limitations in resolving the small physical mass scales involved, we employ high-mass proxy simulations to capture the gas dynamics, followed by a physically motivated renormalization.
Results. We demonstrate that the dynamics of the tidal fallback and subsequent viscous evolution naturally reproduce the fundamental temporal characteristics of observed flares: the peak luminosity and the characteristic 1-hour duration. We show that this fueling mechanism is dynamically viable and energetically consistent, placing strong constraints on the required efficiency of the non-thermal emission process, suggesting extreme radiative inefficiency.
Conclusions. These findings provide compelling evidence for a hidden population of brown dwarfs in the Galactic Center. Crucially, the observed high flare frequency implies tight orbits characteristic of advanced inspirals. This establishes a direct link between electromagnetic transients and active gravitational wave sources, alerting the LISA consortium years in advance to the presence of specific XMRI systems promising exceptionally high signal-to-noise ratios for precision tests of general relativity.
Key Words.:
Galaxy: center – accretion, accretion disks – gravitational waves – stars: low-mass
@@ 1 Introduction
While the emission mechanism is identified as non-thermal synchrotron radiation, the mechanism triggering these events—specifically, the source of the fuel and the trigger for particle acceleration—remains debated. Proposed fueling mechanisms include the interaction with the G2 object, accretion instabilities driven by stellar winds, or large-scale magnetic activity (Pfuhl et al. 2015; Witzel et al. 2014; Ponti et al. 2015; Calderón et al. 2025; Ressler et al. 2019).
We investigate the hypothesis that these flares are fueled by the tidal stripping of material from a brown dwarf (BD) during a grazing encounter with Sgr A*. Stellar dynamics models predict a substantial population of BDs in the Galactic Center. The interaction of these objects with the supermassive black hole has implications beyond electromagnetic flares. BDs inspiraling toward Sgr A* form Extremely-Large Mass Ratio Inspirals (XMRIs), characterized by mass ratios of $\sim 10^{8}$. These systems are key sources of low-frequency gravitational waves detectable by the future Laser Interferometer Space Antenna (LISA) mission (Amaro-Seoane 2019; Amaro Seoane & Zhao 2025). Modeling the electromagnetic signatures of these tidal encounters provides a means to constrain the population of these otherwise undetectable objects.
@@
2 Light curve analysis
We analyze the simulated accretion data to generate synthetic light curves and extract temporal characteristics and variability metrics for comparison with observations. We adopt a methodology that decouples the macroscopic hydrodynamic modeling of the fuel supply from the microscopic plasma physics governing the radiation. The SPH simulations provide the mass accretion rate profile, $\dot{M}_{\text{sim}}(t)$. We convert this into an observed luminosity $L(t)$ through a post-processing pipeline.
First, we address SPH resolution limitations by applying a dimensionless rescaling factor, $E_{\text{strip}}$:
$\dot{M}_{\text{phys}}(t)=E_{\text{strip}}\,\dot{M}_{\text{sim}}(t).$ (1)
We then calculate the non-thermal luminosity in the observed NIR/X-ray bands:
$L_{\text{NT}}(t)=\eta_{\text{NT}}\,\dot{M}_{\text{phys}}(t)\,c^{2}.$ (2)
The present study focuses on the dynamical viability of the fueling mechanism. We do not explicitly model the complex plasma kinetics required to accelerate electrons or generate the observed non-thermal spectrum. Therefore, the non-thermal radiative efficiency $\eta_{\text{NT}}$ serves as a phenomenological parameter encapsulating the efficiency of channeling the available accretion energy specifically into the relativistic electrons responsible for the flare emission.
We interpret the optimized value of $\eta_{\text{NT}}$ as a derived physical requirement on the efficiency of the unmodeled non-thermal processes for this dynamical model to match the observed luminosities, given the constraints on the physically plausible stripped mass $M_{\text{phys}}$.
To model the evolution of the material within the accretion disk, we convolve the luminosity calculated in Eq. (2) with a viscous disk kernel characterized by a timescale $t_{\nu}$. This produces a smooth light curve envelope representing the bulk dynamics. The SPH simulations do not resolve the turbulence responsible for the observed stochastic variability. Therefore, to generate realistic light curves for temporal analysis, we impose a variability structure consistent with observations. We modulate the envelope with a red noise signal characterized by an input PSD slope of 2-2 (consistent with observations) and 30% RMS variability, following the methodology of Timmer & König (1995).
We characterize the temporal profile of the resulting light curve, $L_{\text{NT}}(t)$, using standard metrics for duration and morphology. The primary measure of duration is the Full Width at Half Maximum (FWHM). We determine the FWHM by first identifying the peak luminosity of the event, $L_{\text{peak}}=\max(L_{\text{NT}}(t))$. We then define the half-maximum luminosity level, $L_{\text{HM}}=L_{\text{peak}}/2$. The FWHM duration is the time interval during which $L_{\text{NT}}(t)>L_{\text{HM}}$. We identify the start time $t_{\text{start}}$ and end time $t_{\text{end}}$ by locating the crossings of the light curve with the $L_{\text{HM}}$ level, typically employing linear interpolation between discrete data points. The FWHM duration is then
$\Delta t_{\text{FWHM}}=t_{\text{end}}-t_{\text{start}}.$ (3)
To characterize the asymmetry of the flare profile, we calculate the rise time and the fall time. The rise time is defined as the interval required for the luminosity to increase from 10% to 90% of $L_{\text{peak}}$, and the fall time is the interval for the decrease from 90% to 10% of $L_{\text{peak}}$.
3 Results
Figure 1 illustrates the dynamics of the interaction near the periapsis passage for a representative orbit with $R_{p}=120\,R_{\odot}$ and $a=2560\,R_{\odot}$ (a grazing encounter with $\beta\approx 0.63$). The visualization shows the logarithmic column density of the SPH particles. The strong tidal field of Sgr A* stretches and deforms the brown dwarf structure. Material from the outer envelope is stripped away, forming tidal tails. This stripped material subsequently falls back toward the black hole, fueling the accretion event. We recorded the time-dependent mass accretion rate of this material, $\dot{M}_{\text{simulation}}(t)$.
The physical mass involved in the event ($\approx 10^{-7}\,M_{\odot}$) is comparable to the mass resolution of the SPH simulation ($m_{\text{SPH}}=4.0\times 10^{-8}\,M_{\odot}$), necessitating the renormalization described in the Methodology section.
The optimization yields a best fit characterized by $E_{\text{strip}}=10^{-9}$, $\eta_{\text{NT}}=10^{-8}$, and $t_{\nu}=90$ minutes. This implies a physical accreted mass $M_{\text{phys}}\approx 1.5\times 10^{-7}\,M_{\odot}$, which is physically consistent with the mass expected from a grazing tidal encounter.
Figure 2 displays the resulting synthetic light curve. The model successfully reproduces the key temporal features of observed Sgr A* flares. The peak luminosity reaches $8.1\times 10^{35}\,\text{erg/s}$, falling within the target observational range. The duration of the event (FWHM) is 1.03 hours, consistent with the typical 1-hour duration observed. The characteristic rise and fall times (15.5 min and 23.3 min, respectively) also align well with observations.
4 Discussion
This low efficiency is crucial for consistency with multi-wavelength observations. In standard Radiatively Inefficient Accretion Flow (RIAF) models, the thermal emission (peaking in radio/mm bands) typically dominates the non-thermal emission. While the precise partitioning of energy into non-thermal electrons is uncertain (Yuan et al. 2003), assuming a standard scenario where $\eta_{\text{thermal}}\gg\eta_{\text{NT}}$ (e.g., $\eta_{\text{NT}}$ being 1-2% of the total radiated power) would imply a thermal luminosity $L_{\text{thermal}}\approx 10^{38}$ erg/s. Such high radio luminosities are inconsistent with observations, which show that the NIR flare luminosity is comparable to the millimeter luminosity ($L_{\text{NIRflare}}\sim L_{\text{mm}}$) (GRAVITY Collaboration et al. 2020).
The validation of the tidal fueling mechanism establishes a robust dynamical framework, allowing us to move beyond modeling individual events toward characterizing the entire population. We analyze the potential for sustained, repeated flaring activity, as observed by monitoring campaigns (Dodds-Eden et al. 2009; Witzel et al. 2018). Given the required mass per flare ($\Delta M\approx 1.5\times 10^{-7}\,M_{\odot}$), a $0.01\,M_{\odot}$ BD can sustain approximately 67,000 flares.
This maximum frequency corresponds to a tight orbit ($a\approx 271\,R_{\odot}$) with a periastron near the tidal radius ($R_{p}\approx 75.5\,R_{\odot}$, or $8.3\,R_{g}$), consistent with astrometric localization of flares (Gravity Collaboration et al. 2023). This configuration characterizes an XMRI in an advanced stage of inspiral. The lifetime of the BD in this representative configuration is approximately 46 years.
The ability to identify active XMRI systems via their electromagnetic flares represents a paradigm shift in multi-messenger astronomy. It establishes an early warning system, alerting the LISA consortium years in advance to specific, localized sources. These sources are predicted to generate exceptionally high signal-to-noise ratios (SNRs) (Amaro-Seoane 2019; Amaro Seoane & Zhao 2025), potentially eclipsing even those of supermassive black hole mergers. This foresight is critical, providing the necessary lead time to develop the specialized data analysis pipelines required to extract these complex, dominant signals. The scientific reward is profound: these “golden binaries” will enable measurements of Sgr A*’s fundamental parameters with breathtaking precision—constraining its spin to one part in $10^{10}$ and its mass to within a few solar masses (Vázquez-Aceves et al. 2023, 2025). This unprecedented accuracy transforms Sgr A* into a precision laboratory for strong-field gravity, facilitating definitive tests of the Kerr metric. While enhanced GRAVITY observations can corroborate these predictions, LISA will provide the smoking-gun evidence for this EM-GW synergy.
Acknowledgements.
PAS is indebted with Salvador Ardid for the access to the computational resources which allowed to do the simulations and to James Lombardi for his guidance with the SPH code.
Appendix A Methodology
A.1 Brown dwarf model construction
@@
We constructed realistic initial conditions for the brown dwarf (BD) using the Modules for Experiments in Stellar Astrophysics (MESA) code (Eggleton 1983; Paxton et al. 2011, 2013). We generate a one-dimensional stellar evolution model for a $10^{-2}\,M_{\odot}$ object, evolved for a Hubble time. The resulting stabilized structure is characterized by a logarithmic radius $\log_{10}(R/R_{\odot})=-0.9833$ and a logarithmic effective temperature $\log_{10}(T_{\text{eff}}/\text{K})=2.3447$. The composition is defined by mass fractions of hydrogen $X=0.7597$, helium $Y=0.2402$, and metals $Z=10^{-4}$.
@@
A.2 Hydrodynamic integrator
We performed the hydrodynamic simulations using the StarSmasher SPH code (Rasio 1991; Lombardi 1998; Rasio & Lombardi 1999; Lombardi et al. 1999). StarSmasher employs variational equations of motion to evolve the system. It calculates hydrodynamic forces by determining the pressure of each particle as a function of its local properties via the equation of state. To calculate gravitational interactions, the code implements a direct $N$-body summation method accelerated on NVIDIA graphics processing units (GPUs). StarSmasher utilizes a cubic spline function for the smoothing kernel. To manage shocks and prevent unphysical interparticle penetration, the code employs an artificial viscosity prescription coupled with a Balsara Switch.
The stabilized BD model is placed on specified Keplerian orbits characteristic of XMRIs progenitors around Sgr A*, which is modeled as a $4.3\times 10^{6}\,M_{\odot}$ point mass sink.
A.3 Tidal stripping and mass renormalization
We first estimate the parameters of the tidal encounter. We analyze a scenario involving a brown dwarf (BD) with $M_{\text{BD}}=0.01\,M_{\odot}$ orbiting Sgr A* ($M_{\text{BH}}=4.3\times 10^{6}\,M_{\odot}$) with an illustrative periastron distance $R_{p}=120\,R_{\odot}$. Assuming a standard BD radius $R_{\text{BD}}\approx 0.1\,R_{\odot}$, the tidal radius is
@@ A.3 Tidal stripping and mass renormalization
The penetration factor, $\beta=R_{t}/R_{p}\approx 0.63$, characterizes a grazing encounter ($\beta<1$).
For such encounters, analytical models predict that a fraction of the stellar mass is tidally stripped. We estimate the stripped mass using the formalism of tidal energy deposition. The energy injected into the BD structure relative to its binding energy is given by
@@
Adopting a representative tidal coupling coefficient $T(\beta)\sim 10^{-4}$ appropriate for this penetration factor, we find $\Delta E_{\text{tide}}/E_{\text{bind}}\approx 6.3\times 10^{-6}$. For an $n=1.5$ polytropic structure appropriate for a BD, this yields an estimate of the analytically expected stripped mass:
$\Delta M_{\text{analytical}}\approx{3.8\times 10^{-8}\,M_{\odot}.}$ (6)
We now examine the resolution constraints of our Smoothed Particle Hydrodynamics (SPH) simulations. We model the $0.01\,M_{\odot}$ BD using $N=250,000$ particles. This sets the mass resolution of the simulation, defined by the mass of a single SPH particle, $m_{\text{SPH}}$:
@@
The critical constraint arises when considering the physical mass required to power the observed flare, given the parameters derived from our optimization workflow (see Sect. 2). To reproduce the observed data from Sgr A* with the derived non-thermal radiative efficiency $\eta_{\text{NT}}\approx 10^{-8}$, the required physical accreted mass is $M_{\text{physical}}\approx 1.5\times 10^{-7}\,M_{\odot}$.